OpenStudy (anonymous):

A goat is tethered to a stake in the ground with a 5-m rope. The goat can graze to the full length of the rope a full 360 degrees around the stake. How much area does the goat have in which to graze?

OpenStudy (tennistar):

A = pi r^2 A = 5^2(3.14) A = 78.5 sq m

OpenStudy (tennistar):

A is area

OpenStudy (anonymous):

In other words, you have a circle, the radius is 5m, how do you find the area?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yea...what she said...

OpenStudy (anonymous):

thanks :)